OpenStudy (anonymous):

given dy/dx[f(2x)]=f'(x) and f'(1)=1, find f'(2)

OpenStudy (anonymous):

@Hero and @FibonacciChick666 please help

OpenStudy (anonymous):

@thomaster

OpenStudy (fibonaccichick666):

ok, so can you tell me what dy/dx stands for?

OpenStudy (fibonaccichick666):

also, do you have more information?

OpenStudy (anonymous):

No more info, and dy/dx is the change in y in respect to the change in x

OpenStudy (fibonaccichick666):

ok hmm, this is interesting, I have a theory for the problem, but I can't guarantee it

OpenStudy (fibonaccichick666):

integrate both sides wrt x , what do you get?

OpenStudy (fibonaccichick666):

or WAIT easier

OpenStudy (fibonaccichick666):

can you re-write f'(x)?

OpenStudy (fibonaccichick666):

in general

OpenStudy (anonymous):

I tried thinking that since they said \[\Large \frac{ dy }{ dx }\left[ f(2x) \right]=f \prime (x)\]

OpenStudy (fibonaccichick666):

what did you get?

OpenStudy (anonymous):

Isn't it \[\Large f \prime (2x)=f \prime (x) \]

OpenStudy (fibonaccichick666):

or was the answer wrong?

OpenStudy (fibonaccichick666):

yea, that's what I'm thinking

OpenStudy (anonymous):

ok and here are hte answer choices sorry I should've included those

OpenStudy (anonymous):

1/4 1/2 3/4 3/2 5/2

OpenStudy (fibonaccichick666):

what? that's weird

OpenStudy (anonymous):

exactly I'd assume that there would be 1 as an answer

OpenStudy (fibonaccichick666):

uhm , I have to go but, let's see. we'd need dy/dx f(4)=f'(2)

OpenStudy (fibonaccichick666):

so maybe something where you split upthe four into f(1)+f(1)+f(1) but I'm not sure

OpenStudy (anonymous):

Hmm, ok thanks anyways

OpenStudy (fibonaccichick666):

this doesn't make sense to me, @ganeshie8 , any ideas?

zepdrix (zepdrix):

\[\Large\rm \frac{d}{dx}f(2x)=f'(x)\]Apply chain rule to the left side,\[\Large\rm 2f'(2x)=f'(x)\]Evaluate this at x=1,\[\Large\rm 2f'(2)=f'(1)\]Understand how to wrap it up? :)

OpenStudy (anonymous):

That's so cool!

zepdrix (zepdrix):

Ya it's a neat problem! I remember seeing this one like a month ago or something like that.

OpenStudy (anonymous):

So it'd be 1/2 since the LHS is 1 and you divide by 2 on both sides to just get f'(2)

zepdrix (zepdrix):

Yah looks good!

OpenStudy (anonymous):

@FibonacciChick666 yeahhh

OpenStudy (fibonaccichick666):

oh wow I overcomplicated. Sorry, but yea, I can see that working now, as a side note though, that means that every derivative should have the same value since 2x changes to just 2, so you can't get 1 for f'(1) but anyways, go with what zepdrix said becasue it's a better lead than I had